A man is 41 years old and his son is 9.In how many years will be the father be three times as old as his son?
Answer by scott8148(6628)

f = 41

s = 9

f + x = 3(s + x) ___ 41 + x = 3(9 + x) ___ 41 + x = 27 + 3x ___ 14 = 2x